### v2

# American Sign Language Translation ASLP American Sign Language Production

ASLT
American Sign Language Translation
ASLP
American Sign Language Production
BSL
British Sign Language
BLEU
Bilingual Evaluation Understudy
BLSTM
Bidirectional Long Short-Term Memory
BSLCPT
BSL Corpus T
CNN
Convolutional Neural Network
CRF
Conditional Random Field
CSLR
Continuous Sign Language Recognition
CTC
Connectionist Temporal Classification
C4A
Content4All
DL
Deep Learning
DGS
German Sign Language - Deutsche Gebärdensprache
DSGS
Swiss German Sign Language - Deutschschweizer Gebärdensprache
DTW
Dynamic Time Warping
FC
Fully Connected
FF
Feed Forward
GAN
Generative Adversarial Network
GPU
Graphics Processing Unit
GRU
Gated Recurrent Unit
GSLP
German Sign Language Production
G2PT
Gloss to Pose Transformer
HMM
Hidden Markov Model
HPE
Hand Pose Enhancer
ISL
Irish Sign Language
LSTM
Long Short-Term Memory
MHA
Multi-Headed Attention
MTC
Monocular Total Capture
MSE
Mean Squared Error
MSLP
Multilingual Sign Language Production
MDN
Mixture Density Network
mDGS
meineDGS
mDGS-V
meineDGS-Variants
mDGST
meineDGST
mDGST-H
meineDGST-HARD
mDGST-E
meineDGST-EASY
NMT
Neural Machine Translation
NLP
Natural Language Processing
PHOENIX12
RWTH-PHOENIX-Weather-2012
PHOENIX14
RWTH-PHOENIX-Weather-2014
PHOENIX14T
RWTH-PHOENIX-Weather-2014T
POF
Part Orientation Field
PAF
Part Affinity Field
P2TT
Pose to Text Transformer
RELU
Rectified Linear Units
RNN
Recurrent Neural Network
ROUGE
Recall-Oriented Understudy for Gisting Evaluation
SGD
Stochastic Gradient Descent
SLA
Sign Language Assessment
SLR
Sign Language Recognition
SLT
Sign Language Translation
SLP
Sign Language Production
SMT
Statistical Machine Translation
T2GT
Text to Gloss Transformer
T2PT
Text to Pose Transformer
T2P
Text to Pose
T2G2P
Text to Gloss to Pose
TTS
Text to Speech
WER
Word Error Rate
SignDiff: Diffusion Models for American Sign Language Production
Sen Fang1, Chunyu Sui2, Yanghao Zhou3, Xuedong Zhang4
Hongbin Zhong5, Mingyu Zhao1, Yapeng Tian6, Chen Chen7
1 Rutgers University, 2Columbia University, 3National University of Singapore, 4Victoria University
5Georgia Institute of Technology, 6The University of Texas at Dallas, 7University of Central Florida
https://signdiff.github.io
Abstract
In this paper, we propose a dual-condition diffusion pre-training model named SignDiff that can generate human sign language speakers from a skeleton pose. SignDiff has a novel Frame Reinforcement Network called FR-Net, similar to dense human pose estimation work, which enhances the correspondence between text lexical symbols and sign language dense pose frames, reduces the occurrence of multiple fingers in the diffusion model. In addition, we propose a new method for American Sign Language Production (ASLP), which can generate ASL skeletal pose videos from text input, integrating two new improved modules and a new loss function to improve the accuracy and quality of sign language skeletal posture and enhance the ability of the model to train on large-scale data. We propose the first baseline for ASL production and report the scores of 17.19 and 12.85 on BLEU-4 on the How2Sign dev/test sets. We evaluated our model on the previous mainstream dataset PHOENIX14T, and our method achieved the SOTA results. In addition, our image quality far exceeds all previous results by 10 percentage points in terms of SSIM.
Figure 1: The Process of Sign Language Production: (a) Enter an English text you want to translate, our ASLP method then translates the text into a continuous skeleton pose sequence. (b) Use our FR-Net to predict the generated human body shape. Unlike previous pose2video work, FR-Net learn human body shape as an additional input, rather than just skeletal condition; (c) User can use pose-condition human synthesis model [23] to further generate an final sign language video.
I INTRODUCTION
Sign language is a complex language with various challenges for learners in both learning and expression [53]. In sign language processing (SLP), as shown in Fig. 1, textual vocabulary is transformed into skeleton sequences [47, 45, 72, 24, 5] and modeled into 3D human body videos. Some studies also learn gloss, which is a type of sign language annotation that represents text. Previous Sign Language Production (SLP) approaches combined techniques from neural machine translation (NMT), generative adversarial networks (GANs), pose extraction, and image transformation [54, 72, 44]. However, data processing in sign language is time-consuming and complex due to the lack of standardization and the need for multiple rounds of dataset processing [56, 7, 21]. How2Sign [16] is a recently released multimodal and multi-view continuous American Sign Language (ASL) dataset in 2021, whereas previous sign language production studies mainly utilized the RWTH-PHOENIX-Weather-2014T (PHOENIX14T) German sign language dataset [17]. Due to the complexity of data processing, fewer researchers have yet explored ASL production using deep learning techniques on this dataset. Compared with PHOENIX14T, the duration and vocabulary of How2Sign are approximately 7.53 and 5.74 times larger than the commonly used PHOENIX14T, respectively.
In this paper, we propose a large-scale American Sign Language Production (ASLP) method (text2pose) and a diffusion model [41] for pose-conditioned human body synthesis (pose2video). Firstly, we discard the intermediate conversion stage from text to gloss and directly perform a more efficient translation from spoken language to pose (text2pose), the method will be abbreviated as “Fast-SLP” in the following. Then, we propose a SignDiff model to learn the correspondence between sign language keypoints and real images representations, enabling high-quality rendering of the generated long sequential videos (pose2video).
To accommodate the larger and semantically richer dataset How2Sign, we have developed adaptable modules that optimize mask handling, batch computation, memory utilization, and tensor operations for Fast-SLP Transformers. Additionally, we have designed a novel loss function specifically tailored to the model, which enhances the learning process for longer videos and enables the production of sign language demonstrations of extended durations. Furthermore, we apply SignDiff to create higher-resolution and visually appealing representations of real human signers. SignDiff is the first diffusion model dedicated to producing authentic sign language videos. We developed SignDiff by fine-tuning the widely-used Stable Diffusion 1.5 model [41], training it on carefully processed videos from the How2Sign dataset. Previous GAN-based sign language rendering methods [55, 12, 62, 46, 32] could not meet our demand due to the increasing volume of training data.
In order to adapt the diffusion model to sign language tasks, we propose a Frame Reinforcement Network called FR-Net. Previous sign language rendering approaches, although primarily relying on video frames extracted using the popular human pose estimation algorithm OpenPose [7] as the main input, have also started incorporating DensePose [21] as an auxiliary fine-tuning input (Specifically, DensePose enables precise localization and pose estimation of dynamic individuals by establishing mappings between 2D images and 3D human models). As DensePose [21] is not an inherent conditional input in ControlNet [73], we devised a DensePose-like input and applied it in SignDiff. The results demonstrate SignDiff significantly improved image quality, surpassing previous works by more than 10 percentage points. Furthermore, we evaluated our improved text2pose generative model, which exhibited favorable results on both the How2Sign and PHOENIX14T datasets. We also conducted quantitative and qualitative assessments of SignDiff, and the experiments demonstrated its superior visual quality compared to previous related works [65, 66, 8, 45, 55, 48].
The contributions of this paper can be summarised as:
  • We propose the first larger-scale American sign language pose video generative model, it demonstrates significant improvements in various capabilities, including the generation of long skeletal pose videos.
  • The novel ASLP approach consists of two new modules and a new loss function, enhances training efficiency, and achieves SOTA performance in text2pose task.
  • We design a dual-condition SignDiff model, it introduces a novel DensePose-like novel conditional input by FR-Net. This innovation significantly improves pose2video production speed, enhancing image quality, and reducing the number of required training samples.
II RELATED WORK
II-A Sign Language Production
The development within the field of sign language is not well-balanced, with a concentration of work on Sign Language Recognition (SLR) [13, 20, 29, 26, 10, 30] and Sign Language Translation (SLT) [6, 9, 28, 4]. Various approaches have been explored, including those based on traditional models, neural networks, and model optimization methods. These different techniques have been used in the main research directions. With the development and deepening of the field, there has been increased attention given to Sign Language Production (SLP).
Previously, researchers mainly relied on synthetic animation methods for Sign Language Production (SLP) [11, 27, 51, 37, 36]. However, with the introduction of the Transformer model, many studies have emerged based on neural translation and kinematics. These studies typically focused on PHOENIX14T as a means of exploration and experimentation [45, 47, 24, 48, 49]. While many of these works involve an intermediate step of translating text to gloss, the temporal control of sign language production itself is challenging, leading to potential inaccuracies when translating to gloss.
In our study, we do not consider this intermediate step due to the significantly larger dataset that our model uses, which does not require the use of an intermediate step to compensate for data scarcity. Additionally, we improve the effectiveness of translation by optimizing structural modules.
II-B Rendering of Conditional Input
Conditioning refers to the ability of a generative model to control its output according to our intentions like ControlNet [73]. Previous conditional input Generative Adversarial Networks (GANs) [19] demonstrated good performance in image [25, 40, 66, 76, 38] and video generation [35, 61, 63, 64, 65]. There have also been numerous works focusing on generating human poses conditioned on various factors, including whole body [1, 34, 38, 52, 57, 77, 8], face [14, 31, 71, 60, 70, 74] and hand [33, 58, 69].
One specific application is human-style transfer [43], which involves replacing a person in a video with another person while keeping their actions unchanged. This technique has also been widely used in the context of sign language production [8, 68, 75]. The key lies in extracting keypoints to replicate the movements [8, 62], with tools such as OpenPose, i3D, and DensePose being commonly employed for keypoint extraction [8, 68, 75, 39].
In our work, we propose a novel model called SignDiff, which incorporates a DensePose-like network named Frame Reinforcement Network on top of the ControlNet. FR-Net follows the idea of ‘learning human body shape, not just bone position’ and is a good extension based on ControlNet. This model serves as a valuable addition to the field and bridges the gap between diffusion models and sign language.
Figure 2: (a) For our ASLP/text2pose approach, the training data comprises spoken text extracted from videos and pose information ($x$: text, $p$: keypoints json file, $y$: original video frame). (b) The interaction principle of training data and FR-Net. When using SignDiff, the original frames of the video are used to learn human body shape through FR-Net, help SignDiff realize the function of rendering real human images from visual images of pose key points. $\bm{e}_{1}$ to $\bm{e}_{4}$ represents the output (i.e., dual-conditional intermediate representation, it is the superposition of skeleton and shape). The output of FR-Net can exhibit diverse styles, which are influenced by several factors, including image resolution and fine details.
III METHODOLOGY
In this section, we will first give an overview of the Sign Language Production (SLP) pipeline. Then, we will elaborate on the details of our ASLP method and SignDiff.
III-A Motivation and Design Idea
For larger-scale ASLP, the challenge lies in the fact that the previous German gloss vocabulary was around 1000, while our vocabulary exceeds 10000. This necessitates substantial methodological enhancements to accommodate the expanded data. To address this challenge, we have devised and integrated two additional adaptive modules. Furthermore, to effectively handle the heightened semantic complexity, we have introduced a novel loss function tailored for our model. Subsequently, GANs were insufficient in simulating genuine and complex sign language data, leading to distorted representations. To overcome this, we adopted diffusion models as advanced sign language rendering models. Unlike GANs, these likelihood-based models effectively capture complex natural image distributions without encountering mode-collapse or training instabilities, all while leveraging parameter sharing instead of excessive parameters like Autoregressive models.
III-B Dataset Processing
The How2Sign dataset is the first large-scale multimodal and multiview continuous ASL dataset. It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information. Despite its significance, prior researchers in the field of sign language have not effectively employed this corpus due to the intricacies involved in American Sign Language Production (ASLP). Our objective is to train a Sign Language Pose network to generate sequences of 3D skeleton poses on How2Sign as shown in Fig 2. Initially, we extract 2D joint positions from each video using the OpenPose method [7]. We then incorporate the skeletal model estimation improvements introduced in Zelink et al. [72] to transform the 2D joint positions into their corresponding 3D counterparts. Moreover, we employ an iterative inverse kinematics approach to minimize the 3D pose while preserving consistent bone lengths and correcting any misplaced joints. Finally, following a similar approach proposed in Stoll et al. [54], we normalize the skeletal data and adjust the data to a trainable state. Our scripts and preprocessing tools are available in supplementary materials and will be useful to the sign language community.
III-C Preliminary
Text to Pose. The general workflow is given a sequence of text, $\mathcal{X}$, our goal is to directly convert it into a continuous signing pose sequence, $\hat{\mathcal{Y}}=(y_{1},...,y_{\mathcal{T}})$ with $\mathcal{T}$ frames. This translation can be formulated as a sequence-to-sequence problem, similar to the text-to-gloss translation discussed in Saunders et al. [50]. However, instead of generating an intermediate gloss sequence, we directly generate the pose sequence. To achieve this, we utilize and update the encoder-decoder transformer framework used in the gloss-to-pose translation. In the modified framework, the source tokens are the words in the spoken language sequence, and the target tokens are the corresponding pose representations. The encoder-decoder transformer can be defined as follows:
$f_{t}=E_{T2P}(x_{t}|x_{1:\mathcal{T}})$ (1)
$p_{w+1}=D_{T2P}(p_{w}|p_{1:w-1},f_{1:\mathcal{T}})$ (2)
In this formulation, $f_{t}$ and $p_{w}$ represent the encoded source and target tokens respectively, where $p_{w}$ now represents the pose at timestep $w$. Similar to the text-to-gloss translation, we can compute the pose at each timestep as $\hat{y}_{w}=\operatorname*{argmax}_{i}(p_{w})$ until a specific termination condition. The $argmax_{i}$ operation is applied over all possible poses, selecting the pose with the highest predicted probability at each time step w. The $argmax_{i}$ denotes the value of i where we maximize the function $p_{w}$ (the pose probability at time step w). By directly translating the spoken language sequence to the signing pose sequence, we avoid the need for an intermediate gloss representation. This not only simplifies the translation process but also allows for a more direct mapping between spoken language and sign language.
Figure 3: We condition Stable Diffusion model (SD) via concatenation or by a more general cross-attention mechanism, which is now the main theoretical basis for controlling SD.
Latent Diffusion Models. Diffusion Models are probabilistic models that learn a data distribution $p(r)$ by denoising a normally distributed variable in a step-by-step manner [41]. This process is the reverse of a fixed $T$ length Markov Chain. Diffusion Models can be seen as a sequence of denoising autoencoders $\epsilon_{\theta}(r_{t},t);\,t=1\dots T$, each trained to predict a clean version of the input $r_{t}$, as shown in Fig. 3.
$L_{DM}=\mathbb{E}_{r,\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(r_{t},t)\|_{2}^{2}\Big{]}\,,$ (3)
The objective of the model is to minimize the difference between the denoised output $r_{t}$ and the original input $x$, with $t$ uniformly sampled from $\{1,\dots,T\}$. This native diffusion model will then be used in ControlNet’s base model.
Controllable Pose to Video. ControlNet [73] manipulates the input conditions of neural network blocks to exert control over the overall behavior of the entire neural network. Here, a “network block" refers to a set of neural layers that are put together as a frequently used unit to build neural networks, e.g., “resnet” block, “conv-bn-relu” block, multi-head attention block, transformer block, etc. Taking the example of a 2D feature, a feature map $\bm{x}\in\mathbb{R}^{h\times w\times c}$ where $\{h,w,c\}$ represent the height, width, and channel dimensions respectively. A neural network block $\mathcal{F}(\cdot;\Theta)$ consisting of parameters $\Theta$ is employed to transform $\bm{x}$ into another feature map denoted as $\bm{y}$. This transformation is represented by the equation $\bm{y}=\mathcal{F}(\bm{x};\Theta)$, and its graphical representation is shown in Figure 2-a and Figure 4.
Subsequently, this replica is locked and then cloned, with the clone being trainable. This forms the basis for ControlNet training [42, 25], where our FR-Net processes the inputs of the cloned replica. The neural network blocks are connected by a unique type of convolution layer called “zero convolution”, i.e., $1\times 1$ convolution layer with both weight and bias initialized with zeros. The resulting model trained through How2Sign is our SignDiff for pose2video.
III-D Tensor-Optimization-Based New Text2Pose Method
Batch & Computational Optimization. We propose a novel technique, called “Batch Optimization" to improve the efficiency of mini-batch processing in deep learning architectures. Our approach leverages foundational tensor operations to construct the source mask tensor with appropriate dimensions, reducing the need for extensive tensor manipulations. By directly manipulating the target input tensor, we introduce memory optimization strategies that streamline computations under certain conditions. This avoids the explicit concatenation of tensors within loops and instead utilizes efficient in-place operations. These optimization strategies collectively aim to circumvent memory inefficiency pitfalls, resulting in substantial improvements in memory utilization and computational speed.
In addition, we introduce a mask optimization technique called “Computational Optimization" in code to enhance the efficiency of loss computation. The mask, a binary tensor of the same shape as the target tensor, determines whether each corresponding target value is a padding value or not. Applying this mask to the predictions and targets eliminates the need to calculate the loss for padding values, reducing unnecessary computations. Given a tensor $T$ and the pad_index used to mark the padding values, we can define a mask tensor $M$ using the following formula:
$M_{ij}=\begin{cases}1,&\text{if }T_{ij}\neq\text{pad\_index}\\ 0,&\text{otherwise}\end{cases}$
Here, $i$ and $j$ represent the indices of the tensor $T$. The mask tensor $M$ has the same shape as the target tensor $T$, and each element in the tensor is 1 if the corresponding element in $T$ is not a padding value, and 0 otherwise. Formally, $P$ denotes the tensor of predictions and $T$ denotes the tensor of targets. The masked predictions $P_{masked}$ and masked targets $T_{masked}$ are obtained as $P_{masked}=P\times M$ and $T_{masked}=T\times M$, respectively, using element-wise multiplication ($\times$). The loss is then computed using the selected loss function applied to the masked predictions and targets, such as the default MSE loss or our SA loss.
Scale-Adjusted Loss. It aims to improve the generalization of our model and address the overfitting problem when handling extremely large-scale datasets, which is then calculated using a specific criterion, which takes the $preds_{masked}$ and $targets_{masked}$ mentioned in the previous paragraph as inputs. To adjust the magnitude of the loss, a loss scale factor is applied. If the loss scale is not equal to 1.0, the loss is multiplied by the loss scale factor. Additionally, SA regularization is performed to prevent overfitting. The regularization loss, $SA_{loss}$, is initialized as a tensor with a value. The SA_loss is initialized only before the optimization loop begins. However, SA_loss is updated within the loop by adding the SA norm of each weight parameter containing the string ’weight’. The SA norm of the weight parameters in the model is calculated using the torch.norm function. The computed SA norm is then added to the $SA_{loss}$ variable for each weight parameter that contains the ‘weight’ string in its name. A regularization coefficient, $SA_{lambda}$, is defined with a value of 0.01 to control the strength of the regularization. The SA regularization loss, the result of $SA_{lambda}$ $SA_{loss}$ will be added to the total loss.
III-E SignDiff for Efficient Pose2Video Production
Figure 4: The $\bm{p}_{w}$, $\bm{e}_{j}$ and $\bm{y}_{t}$ represent profound characteristics within neural networks. The term “zero convolution" denotes a convolution layer with dimensions of 1 × 1, where the weight and bias parameters are initialized to zeroes. This diagram shows how we can add a second extra condition to the ControlNet.
Our SignDiff is used to render visual images of actions generated by ASLP into real images with the same actions. The FR-Net is employed to learn the shape of the human body but not only skeletal pose, creating an intermediate representation during rendering, which enhances the accuracy of synchronizing human body movements.
Frame Reinforcement Network For SignDiff’s training, we modified the ControlNet [73] architecture as shown in Figure 4. Specifically, we introduced a new conditional input next to ControlNet for capturing the shapes of people. The frame data collected from ground truth was processed by FR-Net to obtain more specific conditional embeddings. This ensures that the generated results are not distorted solely due to the skeletal structure, but also incorporate the learned aspects of human physique. In addition, Controllable SD can reduce the big data crash problems that GANs may encounter.
FR-Net utilizes adaptive thresholding to create a binary image based on local pixel thresholds. Subsequently, a 5x5 convolutional kernel is utilized for erosion, shrinking white regions in the binary image. This helps refine foreground object segmentation. ControlNet [73] locks all parameters in $\Theta$ and then clones it into a trainable copy $\Theta_{\text{c}}$. The copied $\Theta_{\text{c}}$ is trained with an external condition vector $\bm{c}$. We extend a copy of $\Theta_{\text{d}}$ and define the new input as $\bm{d}$. Here $\bm{c}$ represents OpenPose’s input to the model, and $\bm{d}$ represents our FR-Net’s input to the model. Since our input is ultimately superimposed, we can call the variable that $\bm{c}$ and $\bm{d}$ superimpose $\bm{e}$. Overall, we follow the same principles as ControlNet to incorporate new input conditions.
In principle, as mentioned by the Sec. 3, we denote the zero convolution operation as $\mathcal{Z}(\cdot;\cdot)$ and use two instances of parameters $\{\Theta_{\text{z1}},\Theta_{\text{z2}}\}$ to compose the ControlNet structure with
$\bm{y}_{\text{e}}=\mathcal{F}(\bm{x};\Theta)+\mathcal{Z}(\mathcal{F}(\bm{x}+\mathcal{Z}(\bm{e};\Theta_{\text{z1}});\Theta_{\text{e}});\Theta_{\text{z2}})$ (4)
where $\bm{y}_{\text{e}}$ becomes the output of this neural network block, as visualized in Fig. 4-middle. The experimental results show that our FR-Net can reduce Dynamic Time Warping [^1] DTW [3] is the technique used to dynamically warping the time axis of both sequences based on a defined matching criterion. scores as well as SignGAN [50].
TABLE I: Our DL-Based American Sign Language Pose Video Production Baseline.
DEV SET TEST SET
Type: BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE
Fast-SLP (Ours) 17.19 23.11 29.49 36.96 55.85 12.85 17.35 23.38 39.46 46.89
TABLE II: Fast-SLP Transformer results for Text to Sign Pose production on the How2Sign, with multiple data augmentation techniques. FP: Future Prediction, GN: Gaussian Noise, BO: Batch Optimization, CO: Computational Optimization.
DEV SET TEST SET
Approach: BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE BLEU-4 BLEU-3 BLEU-2 BLEU-1 ROUGE
Base 9.86 13.65 20.99 37.36 40.85 8.84 9.86 17.30 36.88 37.30
Future Prediction 12.67 17.33 25.93 43.68 45.14 12.82 15.05 23.88 41.36 43.24
Just Counter 14.05 20.16 27.71 45.35 47.20 12.80 15.72 25.36 43.76 44.96
Gaussian Noise 15.12 21.92 29.98 47.32 49.74 14.57 18.86 27.03 44.17 46.30
FP & GN 15.18 21.94 29.82 47.13 49.46 15.92 19.65 27.91 46.25 46.57
Scale-Adjusted Loss 14.42 20.84 28.33 44.77 47.08 15.12 18.66 26.51 44.03 44.24
BO & CO 16.92 22.43 30.01 45.15 48.75 12.28 15.99 22.67 37.58 37.84
IV Experiments
In this section, we evaluated our Fast-SLP, SignDiff and FR-Net methods. The experimental content includes back translation assessment, baseline comparison, ablation experiments, and training efficiency tests.
IV-A Experimental Setup
As mentioned above, our training is conducted on datasets such as How2Sign and PHOENIX-14T, and we also evaluate our models separately on the same datasets. The How2Sign training dataset consists of 31,048 video segments, with 1,739 segments in the development set and 2,343 segments in the test set. In our preprocessing stage, we first discard text segments that do not have corresponding videos and then batch-separate them into frames. Further processing is applied for different models. The skeletal data is converted into h5 files, followed by filtering out erroneous data (we read the relevant data and delete the unqualified data such as NaN, 0, or replace it with the average median of the data. Then, we condensed the data to about 1/5 of the original). In the end, we obtain 3D skeleton data in the same format as the base model [45] to train and test our Text2Pose module’s methods. Thousands of files are processed and optimized before being read into the final result.
After preprocessing, the How2Sign dataset contains 31,047, 1,739, and 2,343 segments respectively. Due to computational limitations and cost considerations, as well as ensuring fair evaluations, we train our models on approximately 8,000 segments from the How2Sign train set. As the first baseline in ASLP, although it is unnecessary to compare our work with previous ones, we also conducted tests using our method on previous mainstream datasets: 1) For experiments that need to be compared to previous work, we run their PHOENIX-14T dataset, such as the back-translation evaluation in the case of Germanic languages. 2) For experiments that do not need to be compared to previous work, we train and test on How2Sign, such as the ablation experiments, multiple evaluations of SignDiff, and Qualitative Evaluation. The training/test data for other models will be mentioned in the Experimental section.
Our translation model utilizes the video-level translation model developed by Tarrés et al. [59] and translation model [9], which are refined with retraining. However, due to their model’s lower performance than previous best level of mature sign language interpretation, it fails to fully showcase the accuracy of our modelf in ASL tasks. Thus, we hope that future researchers can further explore and expand in American Sign Language Translation (ASLT).
IV-B Evaluation for Fast-SLP
Back Translation of ASL. Back translation refers to the process of translating generated videos into spoken sentences, which are then compared with the input sentences. This serves as a popular metric for evaluating SLP [45]. BLEU is a commonly used evaluation metric for this method, with higher values indicating a smoother expression of meaning in the videos. We conducted tests using our new method on the How2Sign dataset and proposed a baseline for the first DL-based ASL video production task, as shown in Table I.
TABLE III: Results on the How2Sign dataset for the Text to Pose task, different sign languages have different poses.
DEV SET TEST SET
Approach: BLEU-4 ROUGE BLEU-4 ROUGE
Progressive Transformers [45] 15.18 49.46 15.92 46.57
Neural Sign Actors [2] 13.12 47.55
Fast-SLP Transformers (Ours) 17.19 55.85 12.85 46.89
In Table III, we also used previous method to train and test it on How2Sign ASL datasets for comparison with previous work. We now use one of these jobs as a baseline to run our new data set for comparison. The result shows that our approach is also competitive for excellent existing work, multiple sign languages, not just for data quality reasons. NSA [2], which appears after our ASLP work, shows that their performance is only roughly on par with ours.
Figure 5: Comparison of different settings on DTW values (the lower the better) at different training times.
Ablation Study. As shown in Table II, firstly, we assess the performance of the model on the text-to-pose task described in Section III-C. We evaluate two new modifications to the baseline model and introduce a new loss function. The evaluation is conducted on the development testing set of How2Sign. We examine the effectiveness of the baseline model without any modifications and also explore the results of various settings that provide improvements.
In Table II, it can be observed that Gaussian Noise enhances the sensitivity of the model toward isolated word recognition. The combination of FP and GN further improves the adaptability of the model to unknown sentences. The Scale-Adjusted Loss performs well while emphasizing the temporal characteristics of sign language. This method shows moderate improvements over the base model, particularly in higher-order BLEU scores (BLEU-3 and BLEU-4) on both dev and test sets. Its performance suggests that it effectively captures the sequential nature of sign language, which is crucial for maintaining coherence in longer sign language productions. Finally, we discovered that our batch module is better suited for longer English sentences, as it achieves a commendable performance. This is particularly evident in the dev set results, where it outperforms all other methods in BLEU-4 (16.92) scores. These higher-order BLEU scores are more indicative of the model’s ability to capture longer phrase structures and overall sentence coherence.
(a) Inputs (b) Generated images
Figure 6: We presented a comparison of SignDiff with previous work [73, 65], including (a) input and (b) generated images. We demonstrated the effect of training 10,000 images and 50,000 images, which demonstrates that we only need fewer samples and have higher training efficiency. We also demonstrated the effects that two models of native stable diffusion can achieve for comparison. In reality style, sign language is accurate but prone to missing fingers, while in anime style, fingers are complete but not accurate.
Training Efficiency Study. In Figure 5, we employ DTW histograms to evaluate the performance of our model during training, without the necessity of complete training. The DTW algorithm, based on the concept of dynamic programming, is utilized to calculate the minimum operation distance between sign language videos and sentence sequences. It facilitates the computation of visual-textual matching in the latent space, with lower values indicating better results. Additionally, we overlay another line plot of the loss during training to comprehensively assess the effectiveness of our modifications and the performance of the model, providing guidance for future work.
TABLE IV: Back translation results on the PHOENIX-14T dataset for the Text to Pose task.
DEV SET TEST SET
Approach: BLEU-4 ROUGE BLEU-4 ROUGE
Progressive Transformers [45] 11.82 33.18 10.51 32.46
Adversarial Training [44] 12.65 33.68 10.81 32.74
Mixture Density Networks [47] 11.54 33.40 11.68 33.19
Mixture of Motion Primitives [48] 14.03 37.76 13.30 36.77
Photo-realistic SLP [50] 16.92 35.74 21.10 42.57
Fast-SLP Transformers (Ours) 18.26 39.62 22.15 46.82
Back Translation of GSL. This table IV presents the back translation results for various German Sign Language production methods on the PHOENIX-14T dataset [45, 44, 47, 48] for the “Text to Pose" task. We can observe a clear progression in the field from the earlier Progressive Transformers to our proposed Fast-SLP Transformers. The initial methods such as Progressive Transformers and Adversarial Training achieved relatively lower BLEU-4 and ROUGE scores on both the development and test sets. Subsequently, approaches like Mixture Density Networks and Mixture of Motion Primitives showed improvements in performance. Notably, the Mixture of Motion Primitives method demonstrated significant advancements over its predecessors, they made many useful attempts in accurate pose prediction.
The most recent two approaches, Photo-realistic SLP and our proposed Fast-SLP Transformers, showcase the latest advancements in the field. Photo-realistic SLP achieved a notable performance leap on the test set, particularly in terms of BLEU-4 scores. However, our Fast-SLP Transformers method outperformed all previous work across all metrics, achieving the highest BLEU-4 and ROUGE scores on both the development and test sets. This indicates that our approach has made substantial progress in generating more accurate and fluent sign language translations, setting a new benchmark for future research in this area.
IV-C Evaluation for SignDiff
Finally, we evaluated the image generation performance of SignDiff, including both qualitative and quantitative evaluations (baseline comparison and ablation and study).
Qualitative Evaluations. Figure 6 showcases three distinct input types and the corresponding generation results from different models. A comparison between our model and Vid2Vid reveals that our approach achieves superior results with less data input. Upon convergence, our model surpasses the performance of Vid2Vid in terms of generation quality. Moreover, when contrasted with a common LoRA [23] model used to generate human figures, our approach demonstrates clear advantages. While stable diffusion models excel in producing accurate gestures, they often lack detail in finger representation. Conversely, the cartoon style preserves finger details but sacrifices some precision in overall gesture depiction. Our model successfully combines the strengths of both approaches, offering both gesture accuracy and exceptional image clarity. Furthermore, as discussed in Section III-A, our method demonstrates superior image quality upper bounds and enhanced stability. This comprehensive improvement in various aspects of image generation underscores the effectiveness of our approach in addressing the challenges of sign language synthesis.
Baseline Comparison. We first compared it with previous advanced methods [8, 65, 66, 55, 50], where the input is the corresponding conditional input. We compared the output with Ground Truth and calculated the indicator of image quality as follows: 1) Structural Similarity (SSIM) was computed over the entire image [67]. 2) Hand SSIM was calculated by applying the SSIM metric to a cropped region of each hand. 3) Hand Pose was determined by measuring the absolute distance between the 2D hand keypoints in the generated and ground truth hand images, utilizing a pre-trained hand pose estimation model [18]. 4) Fréchet Inception Distance (FID) was evaluated for the complete image [22].
TABLE V: Comparison of Sign Language Video Generation Results: SignDiff vs. Existing Models.
SSIM $\uparrow$ Hand SSIM $\uparrow$ Hand Pose $\downarrow$ FID $\downarrow$
EDN [8] 0.737 0.553 23.09 41.54
vid2vid [65] 0.750 0.570 22.51 56.17
Pix2PixHD [66] 0.737 0.553 23.06 42.57
Stoll et al. [55] 0.727 0.533 23.17 64.01
SignGAN [50] 0.759 0.605 22.05 27.75
SignDiff (Ours) 0.849 0.676 20.04 25.22
Table V demonstrates that our SignDiff significantly outperforms existing state-of-the-art approaches across all evaluation metrics. We achieve substantial improvements in image similarity, with an SSIM score of 0.849 and Hand SSIM of 0.676, surpassing previous benchmarks. Our method also shows superior performance in Hand Pose accuracy (20.04) and perceptual quality (FID score of 25.22). These improvements can be attributed to the strengths of our diffusion model architecture and targeted optimizations using high-quality datasets. The FR-Net plays a crucial role by incorporating detailed hand shape information during training, unlike previous approaches that focus primarily on skeletal structures. In conjunction with Table VI, these results highlight the effectiveness of our methodological innovations in advancing sign language video generation. Our approach sets a new standard for quality and accuracy in this important field of assistive technology.
TABLE VI: Comparison of our FR-Net and other approaches that can enhance image quality [50].
SSIM $\uparrow$ Hand SSIM $\uparrow$ Hand Pose $\downarrow$ FID $\downarrow$
Baseline 0.743 0.582 22.87 39.33
Hand Discriminator 0.738 0.565 22.81 39.22
Hand Keypoint Loss 0.759 0.605 22.05 27.75
SignDiff (No FR-Net) 0.817 0.646 21.09 26.02
FR-Net (Ours) 0.849 0.676 20.04 25.22
Figure 7: Comparison of training efficiency between FR-Net and SignDiff (No FR-Net), comparison of the DTW and loss value of different periods (according to the number of pictures trained), the smaller the DTW score, the better.
Ablation Study. We conducted an ablation study on SignDiff to evaluate the impact of various components, particularly the FR-Net. Table VI presents a comparison of our method against several baselines and intermediate approaches. The results demonstrate clear performance improvements across all metrics as we incorporate more sophisticated techniques. Our proposed FR-Net consistently outperforms other methods, showing substantial enhancements in image quality and hand pose estimation. Compared to the baseline, our method achieves significant improvements: a 10.6% increase in overall SSIM, a 9.4% improvement in Hand SSIM, and a 2.83% reduction in Hand Pose error. Notably, the FID score shows a dramatic 14.11% reduction from the baseline to our full model, highlighting the effectiveness of our approach in generating more realistic sign language images. As mentioned in Section I, these improvements address the challenge of motion blur in fast-moving sign language images.
We also assessed the training efficiency of the FR-Net, as shown in Figure 7. While the rate of decrease in loss remains consistent, we observe a marked improvement in the Dynamic Time Warping (DTW) metric, suggesting enhanced capture of temporal aspects in sign language gestures without significantly increasing computational overhead. In summary, our ablation study validates the effectiveness of the FR-Net, demonstrating its crucial role in advancing sign language image enhancement and its potential for improving sign language recognition systems.
V CONCLUSION
Providing high-quality sign language to both learning and user communities is of crucial importance for addressing the shortcomings in the community. In this article, we propose the first large-scale production model for ASL videos and the first pose-to-video model based on diffusion models. We introduce a novel FR-Net for the deep integration of sign language and diffusion models. Additionally, we propose two new modules and a new loss function to adapt to our further developed How2Sign ASL dataset on a larger scale. Through baseline comparisons, ablation studies, various parameter analyses, and comprehensive figures, as well as qualitative evaluations, we demonstrate the effectiveness of our approach. In future work, we will enhance the end-to-end capabilities of the model, enabling our technology to be applied to large-scale sign language production.
ETHICAL IMPACT STATEMENT
1. Did you read the Ethical Impact Statement Guidelines document?
Yes, we have read the guidance document on the declaration of ethical influence.
2. Is it clear that all studies and procedures described in the paper were approved (or exempted) by a valid ethical review board? Alternatively, is a valid and sufficient justification provided for why the oversight of an ethical review board was not required?
The data we use is all public data, following the guidelines and licenses provided by the providers. Therefore, for the data itself, we do not assume responsibility, but only as users.
3. Does the ethical impact statement provide a clear, complete, and balanced discussion of the potential risks of individual harm and negative societal impacts associated with the research?
Our SLP model may contain errors or inaccuracies. Users should exercise caution and verify signs before important use [15]. We are not responsible for misinterpretations or misunderstandings.
4. Does the ethical impact statement describe reasonable, valid, and sufficient use of risk-mitigation strategies by the authors to lessen these potential risks? Alternatively, if relevant strategies were not used, is a valid and sufficient justification for this provided?
Sign language is nuanced, requiring users to confirm accuracy. We strive to improve our technology but cannot guarantee error-free results. Users must take necessary precautions for clear and accurate communication.
5. Does the ethical impact statement provide a valid and sufficient justification for how/why the potential risks of the research are outweighed by the risk-mitigation strategies and potential benefits of the research?
Our model has strong features tailored for sign language, which will be notably beneficial for the sign language community and researchers. As it stands, it has not reached a fully commercial level yet. Professional signers can quickly discern errors, and novice signers will not rely on it for learning. In a word, continuing research on SLP technology is currently more beneficial than harmful.
6a. Does the main paper describe whether/how informed consent and/or assent were obtained from participants?
Not applicable, we follow the license to use public data.
6b. Does the main paper state whether the participants explicitly consented to the use of their data in the manner described in the paper?
Not applicable, we follow the license to use public data.
6c. Does the main paper explain whether/how participants were compensated?
Not applicable, we follow the license to use public data.
6d. If the research involves any special or vulnerable populations?
Not applicable, we follow the license to use public data.

### v3

@@
 Word Error Rate
-SignDiff: Diffusion Models for American Sign Language Production
+SignDiff: Diffusion Model for American Sign Language Production
 Sen Fang1, Chunyu Sui2, Yanghao Zhou3, Xuedong Zhang4
-Hongbin Zhong5, Mingyu Zhao1, Yapeng Tian6, Chen Chen7
+Hongbin Zhong5, Yapeng Tian6, Chen Chen7
 1 Rutgers University, 2Columbia University, 3National University of Singapore, 4Victoria University
@@
 In addition, our image quality far exceeds all previous results by 10 percentage points in terms of SSIM.
-Figure 1: The Process of Sign Language Production: (a) Enter an English text you want to translate, our ASLP method then translates the text into a continuous skeleton pose sequence.
-(b) Use our FR-Net to predict the generated human body shape.
-Unlike previous pose2video work, FR-Net learn human body shape as an additional input, rather than just skeletal condition; (c) User can use pose-condition human synthesis model [23] to further generate an final sign language video.
+Figure 1: Our Sign Language Production: First, input the English text you want to translate, and our new Text2Pose method will translate the text into continuous skeletal pose sequences.
+Then we plot the pose data of each frame into appropriate images, and users can further generate the final video using our new pose-conditioned human synthesis model (SignDiff).
+Collaborator Author.
 I INTRODUCTION
@@
 II-B Rendering of Conditional Input
-Conditioning refers to the ability of a generative model to control its output according to our intentions like ControlNet [73].
-Previous conditional input Generative Adversarial Networks (GANs) [19] demonstrated good performance in image [25, 40, 66, 76, 38] and video generation [35, 61, 63, 64, 65].
+Conditioning refers to the ability to control its output according to our intentions like ControlNet [71].
+Previous conditional input Generative Adversarial Networks (GANs) [19] showed good performance in image [24, 39, 64, 74, 37] and video generation [34, 59, 61, 62, 63].
 There have also been numerous works focusing on generating human poses conditioned on various factors, including whole body [1, 33, 37, 51, 56, 75, 8], face [14, 30, 69, 58, 68, 72] and hand [32, 57, 67].
@@
 [53], we normalize the skeletal data and adjust the data to a trainable state.
-Our scripts and preprocessing tools are available in supplementary materials and will be useful to the sign language community.
 III-C Preliminary
@@
 This not only simplifies the translation process but also allows for a more direct mapping between spoken language and sign language.
-Figure 3: We condition Stable Diffusion model (SD) via concatenation or by a more general cross-attention mechanism, which is now the main theoretical basis for controlling SD.
+Figure 3: We condition Stable Diffusion (SD) via concatenation or by a more general cross-attention mechanism, which is now the main theoretical basis for controlling SD.
+$e_{j}$ is the combined skeleton pose and predicted shape representation, how to get it will be covered in detail in Sec.
+IV-B.
 Latent Diffusion Models.
@@
 The resulting model trained through How2Sign is our SignDiff for pose2video.
-III-D Tensor-Optimization-Based New Text2Pose Method
+IV IMPLEMENTATION DETAILS
+IV-A Tensor-Optimization-Based New Text2Pose Method
 Batch & Computational Optimization.
@@
 If the loss scale is not equal to 1.0, the loss is multiplied by the loss scale factor.
-Additionally, SA regularization is performed to prevent overfitting.
+Additionally, $SA_{loss}$ use regularization to prevent overfitting.
 The regularization loss, $SA_{loss}$, is initialized as a tensor with a value.
@@
 The SA regularization loss, the result of $SA_{lambda}$ $SA_{loss}$ will be added to the total loss.
-III-E SignDiff for Efficient Pose2Video Production
+IV-B SignDiff for Efficient Pose2Video Production
 Figure 4: The $\bm{p}_{w}$, $\bm{e}_{j}$ and $\bm{y}_{t}$ represent profound characteristics within neural networks.
@@
 Here $\bm{c}$ represents OpenPose’s input to the model, and $\bm{d}$ represents our FR-Net’s input to the model.
-Since our input is ultimately superimposed, we can call the variable that $\bm{c}$ and $\bm{d}$ superimpose $\bm{e}$.
+Since our input is ultimately superimposed, we can call the variable that $\bm{c}$ and $\bm{d}$ superimpose $\bm{e}$, as shown in Fig.
+3.
 Overall, we follow the same principles as ControlNet to incorporate new input conditions.
@@
 TABLE II: Fast-SLP Transformer results for Text to Sign Pose production on the How2Sign, with multiple data augmentation techniques.
+Lines 2 show the techniques from previous work [44], and the last two lines are our enhanced methods.
 FP: Future Prediction, GN: Gaussian Noise, BO: Batch Optimization, CO: Computational Optimization.
@@
 Base 9.86 13.65 20.99 37.36 40.85 8.84 9.86 17.30 36.88 37.30
-Future Prediction 12.67 17.33 25.93 43.68 45.14 12.82 15.05 23.88 41.36 43.24
-Just Counter 14.05 20.16 27.71 45.35 47.20 12.80 15.72 25.36 43.76 44.96
-Gaussian Noise 15.12 21.92 29.98 47.32 49.74 14.57 18.86 27.03 44.17 46.30
 FP & GN 15.18 21.94 29.82 47.13 49.46 15.92 19.65 27.91 46.25 46.57
-Scale-Adjusted Loss 14.42 20.84 28.33 44.77 47.08 15.12 18.66 26.51 44.03 44.24
-BO & CO 16.92 22.43 30.01 45.15 48.75 12.28 15.99 22.67 37.58 37.84
-IV Experiments
+Scale-Adjusted Loss (Ours) 14.42 20.84 28.33 44.77 47.08 15.12 18.66 26.51 44.03 44.24
+BO & CO (Ours) 16.92 22.43 30.01 45.15 48.75 12.28 15.99 22.67 37.58 37.84
+V EXPERIMENTS
 In this section, we evaluated our Fast-SLP, SignDiff and FR-Net methods.
 The experimental content includes back translation assessment, baseline comparison, ablation experiments, and training efficiency tests.
-IV-A Experimental Setup
+V-A Experimental Setup
 As mentioned above, our training is conducted on datasets such as How2Sign and PHOENIX-14T, and we also evaluate our models separately on the same datasets.
@@
 The training/test data for other models will be mentioned in the Experimental section.
-Our translation model utilizes the video-level translation model developed by Tarrés et al.
-[59] and translation model [9], which are refined with retraining.
-However, due to their model’s lower performance than previous best level of mature sign language interpretation, it fails to fully showcase the accuracy of our modelf in ASL tasks.
-Thus, we hope that future researchers can further explore and expand in American Sign Language Translation (ASLT).
-IV-B Evaluation for Fast-SLP
+V-B Evaluation for Fast-SLP
 Back Translation of ASL.
@@
 We now use one of these jobs as a baseline to run our new data set for comparison.
-The result shows that our approach is also competitive for excellent existing work, multiple sign languages, not just for data quality reasons.
-NSA [2], which appears after our ASLP work, shows that their performance is only roughly on par with ours.
+It shows that our approach is also competitive for excellent latest work, multiple sign languages, not just for data quality reasons.
+NSA [2], which appears after our ASLP work, shows that their performance is only roughly on par with ours (it just 0.27 more).
 Figure 5: Comparison of different settings on DTW values (the lower the better) at different training times.
@@
 These higher-order BLEU scores are more indicative of the model’s ability to capture longer phrase structures and overall sentence coherence.
-(a) Inputs (b) Generated images
-Figure 6: We presented a comparison of SignDiff with previous work [73, 65], including (a) input and (b) generated images.
-We demonstrated the effect of training 10,000 images and 50,000 images, which demonstrates that we only need fewer samples and have higher training efficiency.
-We also demonstrated the effects that two models of native stable diffusion can achieve for comparison.
-In reality style, sign language is accurate but prone to missing fingers, while in anime style, fingers are complete but not accurate.
+Figure 6: Our work is divided into two parts: text2pose and pose2video.
+In this figure, we provide qualitative images from Fast-SLP Transformers and SignDiff respectively.
+(a) shows a comparison between predicted pose images and actual pose images, (b) compares our SignDiff rendered images with the original frames.
+SignDiff uses 2D skeleton visualization images as input to ensure fairness, allowing us to focus on comparing our reconstruction quality.
 Training Efficiency Study.
@@
 Photo-realistic SLP achieved a notable performance leap on the test set, particularly in terms of BLEU-4 scores.
-However, our Fast-SLP Transformers method outperformed all previous work across all metrics, achieving the highest BLEU-4 and ROUGE scores on both the development and test sets.
-This indicates that our approach has made substantial progress in generating more accurate and fluent sign language translations, setting a new benchmark for future research in this area.
-IV-C Evaluation for SignDiff
+However, our Fast-SLP Transformers method outperformed all previous work in all metrics, achieving the highest BLEU-4 and ROUGE scores in both development and test sets.
+The performance of our text2pose model can be directly observed Figure 6 (a), which indicates that our approach has made substantial progress in generating more accurate and fluent sign language translations, setting a new benchmark for future research in this area.
+V-C Evaluation for SignDiff
 Finally, we evaluated the image generation performance of SignDiff, including both qualitative and quantitative evaluations (baseline comparison and ablation and study).
 Qualitative Evaluations.
-Figure 6 showcases three distinct input types and the corresponding generation results from different models.
-A comparison between our model and Vid2Vid reveals that our approach achieves superior results with less data input.
-Upon convergence, our model surpasses the performance of Vid2Vid in terms of generation quality.
-Moreover, when contrasted with a common LoRA [23] model used to generate human figures, our approach demonstrates clear advantages.
-While stable diffusion models excel in producing accurate gestures, they often lack detail in finger representation.
-Conversely, the cartoon style preserves finger details but sacrifices some precision in overall gesture depiction.
-Our model successfully combines the strengths of both approaches, offering both gesture accuracy and exceptional image clarity.
-Furthermore, as discussed in Section III-A, our method demonstrates superior image quality upper bounds and enhanced stability.
-This comprehensive improvement in various aspects of image generation underscores the effectiveness of our approach in addressing the challenges of sign language synthesis.
+Figure 6 (b) shows the quality comparison between our SignDiff reconstructed videos and the original videos.
+To ensure fairness, we used the same 2D pose visualization images as input, and our reconstruction quality is almost identical to the original.
+Furthermore, the diffusion model does not produce motion blur, with each frame showing very clear finger details.
+Our only limitation, as shown in the sixth column, is that hand orientations can become confused, but this minor flaw does not overshadow the overall achievements.
 Baseline Comparison.
@@
 In summary, our ablation study validates the effectiveness of the FR-Net, demonstrating its crucial role in advancing sign language image enhancement and its potential for improving sign language recognition systems.
-V CONCLUSION
+VI DISCUSSION
+We note a concurrent work [2] published before us that adopts an SMPL-based technical approach rather than skeletal poses.
+Both approaches have their advantages and disadvantages.
+While the former may have advantages in pure motion generation, for sign language which requires more user-centric approaches, skeletal poses currently can render more realistic and appealing videos for users.
+VII CONCLUSION
 Providing high-quality sign language to both learning and user communities is of crucial importance for addressing the shortcomings in the community.

### v4

@@
 In addition, we propose a new method for American Sign Language Production (ASLP), which can generate ASL skeletal pose videos from text input, integrating two new improved modules and a new loss function to improve the accuracy and quality of sign language skeletal posture and enhance the ability of the model to train on large-scale data.
-We propose the first baseline for ASL production and report the scores of 17.19 and 12.85 on BLEU-4 on the How2Sign dev/test sets.
+We propose a simple baseline for ASL production and report the scores of 17.19 and 12.85 on BLEU-4 on the How2Sign dev/test sets.
 We evaluated our model on the previous mainstream dataset PHOENIX14T, and our method achieved the SOTA results.
@@
 Furthermore, we apply SignDiff to create higher-resolution and visually appealing representations of real human signers.
-SignDiff is the first diffusion model dedicated to producing authentic sign language videos.
+SignDiff is the a diffusion model dedicated to producing authentic sign language videos (pose2video).
 We developed SignDiff by fine-tuning the widely-used Stable Diffusion 1.5 model [42], training it on carefully processed videos from the How2Sign dataset.
@@
 The contributions of this paper can be summarised as:
-  • We propose the first larger-scale American sign language pose video generative model, it demonstrates significant improvements in various capabilities, including the generation of long skeletal pose videos.
+  • We propose a larger-scale American sign language pose video generative model, it demonstrates significant improvements in various capabilities, including the generation of long skeletal pose videos.
   • The novel ASLP approach consists of two new modules and a new loss function, enhances training efficiency, and achieves SOTA performance in text2pose task.
@@
 II-A Sign Language Production
-The development within the field of sign language is not well-balanced, with a concentration of work on Sign Language Recognition (SLR) [13, 20, 28, 25, 10, 29] and Sign Language Translation (SLT) [6, 9, 27, 4].
+The development within the field of sign language is not well-balanced, with a concentration of work on Sign Language Recognition (SLR) [13, 22, 30, 27, 10, 31, 17] and Sign Language Translation (SLT) [6, 9, 29, 4].
 Various approaches have been explored, including those based on traditional models, neural networks, and model optimization methods.
@@
 With the development and deepening of the field, there has been increased attention given to Sign Language Production (SLP).
-Previously, researchers mainly relied on synthetic animation methods for Sign Language Production (SLP) [11, 26, 50, 36, 35].
+Previously, researchers mainly relied on synthetic animation methods for Sign Language Production (SLP) [11, 28, 52, 38, 37, 18].
 However, with the introduction of the Transformer model, many studies have emerged based on neural translation and kinematics.
@@
 III-B Dataset Processing
-The How2Sign dataset is the first large-scale multimodal and multiview continuous ASL dataset.
+The How2Sign dataset is a large-scale multimodal and multiview continuous ASL dataset.
 It includes a parallel corpus comprising over 80 hours of sign language videos and corresponding modalities such as speech, English transcripts, and depth information.
@@
 Due to computational limitations and cost considerations, as well as ensuring fair evaluations, we train our models on approximately 8,000 segments from the How2Sign train set.
-As the first baseline in ASLP, although it is unnecessary to compare our work with previous ones, we also conducted tests using our method on previous mainstream datasets: 1) For experiments that need to be compared to previous work, we run their PHOENIX-14T dataset, such as the back-translation evaluation in the case of Germanic languages.
+As a simple baseline in ASLP, although it is unnecessary to compare our work with previous ones, we also conducted tests using our method on previous mainstream datasets: 1) For experiments that need to be compared to previous work, we run their PHOENIX-14T dataset, such as the back-translation evaluation in the case of Germanic languages.
 2) For experiments that do not need to be compared to previous work, we train and test on How2Sign, such as the ablation experiments, multiple evaluations of SignDiff, and Qualitative Evaluation.
@@
 BLEU is a commonly used evaluation metric for this method, with higher values indicating a smoother expression of meaning in the videos.
-We conducted tests using our new method on the How2Sign dataset and proposed a baseline for the first DL-based ASL video production task, as shown in Table I.
+We conducted tests using our new method on the How2Sign dataset and proposed a baseline for the DL-based ASL video production task, as shown in Table I.
 TABLE III: Results on the How2Sign dataset for the Text to Pose task, different sign languages have different poses.
@@
 Providing high-quality sign language to both learning and user communities is of crucial importance for addressing the shortcomings in the community.
-In this article, we propose the first large-scale production model for ASL videos and the first pose-to-video model based on diffusion models.
+In this article, we propose a large-scale production model for ASL videos and a pose-to-video model based on diffusion models.
 We introduce a novel FR-Net for the deep integration of sign language and diffusion models.